The force system shown consist of a Couple C and the 4 Forces. The resultant of this system is 500kNm counter-clockwise couple. a. What is the value of P? b. What is the value of Q? c. What is the value of C?


1
Expert's answer
2021-10-14T18:13:57-0400




"R_x=\sum_iF_{xi}=0=>-\dfrac{12}{13}Q+\dfrac{4}{5}P+80kN=0"

"R_y=\sum_iF_{yi}=0=>-\dfrac{5}{13}Q+\dfrac{3}{5}P-20kN=0"

Solving these equations simultaneously gives


"-\dfrac{12}{13}Q+\dfrac{4}{5}P+80kN=0"

"-\dfrac{3}{13}Q+\dfrac{1}{5}P=\dfrac{5}{13}Q-\dfrac{3}{5}P"

"Q=\dfrac{13}{10}P"

"-\dfrac{6}{5}P+\dfrac{4}{5}P+80kN=0"

"P=200kN, Q=260kN"

Given that "C^{R}=500kN\cdot m" counterclockwise, and choosing point "A" as the moment center, we have

"C^{R}=\sum_iM_{Ai}"

"=>500kN\cdot m=-20kN(3m)-C+80kN(4m)"

"+\dfrac{3}{5}P(6m)+\dfrac{4}{5}P(6m)"

Substitute "P=200kN"


"C=-500kN\cdot m-60kN\cdot m+320kN\cdot m"

"+1920kN\cdot m"

"C=1680kN\cdot m"

Because the values for "P, Q," and "C" are positive, each force acts in the direction shown in the figure.


"P=200kN, Q=260kN"

"C=1680kN\cdot m"